OpenStudy (anonymous):

Dada la función f(x)=x^4-2x^2 Responder: ¿Cuales son los Puntos críticos? ¿Intervalos de crecimiento y decrecimiento? Hallar los Puntos de inflexión Calcular los Intervalos de concavidad Hacer un bosquejo de la gráfica

OpenStudy (anonymous):

spanish??

Parth (parthkohli):

1) Find f'(x) and equate that to 0. 2) Graph and see. 3) Inflection points, Google them. 4) ?? 5) Graph?

OpenStudy (turingtest):

Translation: given the function f(x)=x^4-2x^2 what are the critical points? what are the intervals of increasing and decreasing? locate the inflection points find the intervals of concavity make a sketch of the graph

OpenStudy (anonymous):

f(x)=x^4-2x^2 f'(x)=4x^3-4x f'(x)=0=4x^3-4x f'(x)=0=4x(x^2-1) x=0,x=-1,x=1 criticos valores f(0)=0, f(1)=-1.,f(-1)=-1 ....(0,0),(1,-1),(-1,-1) puntos criticos x<-1 decrecimiento -1<x<0 crecimiento 0<x<1 decrecimiento x>1 crecimiento f''(x)=0=12x^2-4=0 x=-0.58 and x=0.58 f(-0.58)=-0.56 f(0.58)=-0.56 (-0.58,-0.56), (0.58,-0.56) Puntos de inflexión

OpenStudy (anonymous):

ok thnx @TuringTest

OpenStudy (anonymous):

\[IN THE GRAPH???\]

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

\[F(X)=X ^{4}-2X ^{2}\]

OpenStudy (anonymous):